OpenStudy (tabithax3):

how do you find f o f if f(x)=1/x+1

OpenStudy (harsimran_hs4):

f o f (x) = f(f(x)) can you figure out now?

OpenStudy (tabithax3):

@harsimran_hs4 i know it's \[f \left( \frac{ 1 }{ x+1 } \right)=\frac{ 1 }{ \frac{ 1 }{ x+1 }+1 } \] but i don't know what to do after

OpenStudy (dumbcow):

simplify ... the easiest yet most difficult part :|

OpenStudy (dumbcow):

first combine denominator into single fraction change 1 to (x+1)/(x+1)

OpenStudy (tabithax3):

@dumbcow so it'll be \[\frac{ 1 }{x+1 }+\frac{ x+1 }{ x+1 }\]

OpenStudy (dumbcow):

yes

OpenStudy (harsimran_hs4):

shouldn`t it be (x+1)/(x+2)

OpenStudy (dumbcow):

haha we're getting there

OpenStudy (tabithax3):

@harsimran_hs4 can you explain for me please?

OpenStudy (tabithax3):

@dumbcow i stopped there, and went blank. help? haha

OpenStudy (dumbcow):

what they posted is the final answer...keep going and you'l get there combine the 2 fractions ... they have a common denominator

OpenStudy (harsimran_hs4):

\[\frac{ 1 }{ \frac{ 1 }{ x + 1} + 1} = \frac{ x+1}{ (x + 1) +1}\]

OpenStudy (harsimran_hs4):

howz it now did you get hold of the final answer

OpenStudy (tabithax3):

@harsimran_hs4 how did you get x+1 on the numerator?

OpenStudy (tabithax3):

oh never mind! got it

OpenStudy (tabithax3):

@harsimran_hs4 i got \[\frac{ x+1 }{ x+2 }\]

OpenStudy (dumbcow):

that was horrible explanation ... you skipped all the intermediate steps

OpenStudy (harsimran_hs4):

cool !!

OpenStudy (dumbcow):

so @tabithax3 , you understand it now

OpenStudy (harsimran_hs4):

finally you got it right!!! @dumbcow i didn`t wanted to go that deep because it would have been spoonfeeding then and i wanted that @tabithax3 should figure that out herself

OpenStudy (tabithax3):

@dumbcow explain? and @harsimran_hs4 but yet my book says the answer should have been \[\frac{ x }{ 2x+1 }\] where \[x \neq-1, -\frac{ 1 }{ 2 }\]

OpenStudy (harsimran_hs4):

ok i get it well we were solving considering f(x) = 1/(1+x) but it was f(x) = (1/x) + 1 this gives the required answer

OpenStudy (tabithax3):

i think i got it, thanks for the help!

OpenStudy (tabithax3):

@harsimran_hs4 i feel stupid now, i wrote the problem wrong all along. oops...but i got the answer right though. sorry for the trouble, but honestly, you really helped me out, and now i understand what to do with the section i'm working on :)

OpenStudy (harsimran_hs4):

happens!!!

OpenStudy (harsimran_hs4):

Welcome